Please, please help! i am stuck on my statistics homework and could use the help to see and know how this is done for future assignments.
here is the assignment:
Resource: Statistics for Psychology
Prepare a written response to the following assignments located in the text:
• Ch. 1, Practice Problems: 12, 15, 19, 20, 21, and 22.
Note. Methods of computation may include the usage of Excel, SPSS™, Lotus®, SAS®, Minitab®, or by-hand computation.

here is one of the questions:
15. Following are the speeds of 40 cars clocked by radar on a particular road in a
35-mph zone on a particular afternoon:
30, 36, 42, 36, 30, 52, 36, 34, 36, 33, 30, 32, 35, 32, 37, 34, 36, 31, 35, 20,
24, 46, 23, 31, 32, 45, 34, 37, 28, 40, 34, 38, 40, 52, 31, 33, 15, 27, 36, 40
Make (a) a frequency table and (b) a histogram. Then (c) describe the general
shape of the distribution.

thank you so much in advance!
bridget

Answer by Edwin McCravy(20060)

The largest value is 52
The smallest value is 15

We will make 6 classes:

To determine the class width
1. Subtract 52-15 = 37
2. Divide by 6. 37÷6 = 6.166666667
3. Round to 6

The smallest class begins at the smallest number

1st class goes from 15 to 15+6 = 21
2nd class goes from 22 to 22+6 = 28 
3rd class goes from 29 to 29+6 = 35
4th class goes from 36 to 36+6 = 42
5th class goes from 43 to 43+6 = 49
6th class goes from 50 to 50+6 = 56

Classes       frequencies
15 - 21           2
22 - 28           4
29 - 35          17 
36 - 42          13
43 - 49           2   
50 - 56           2
----------------------
Total            40

------------------------------------
Make the histogram, starting and ending
the bars halfway between the class bounds.
I put the frequencies on top of the bars.
(Your teacher may vary from this, as there
is no standard way to do this. Some people
put a vertical frequency axis.)



To determine if it is bell-shaped, we calculate the
mean as 34.325 and the standard deviation as 7.325

1 standard deviation below the mean is 34.325 - 7.325 = 27
1 standard deviation above the mean is 34.325 + 7.325 = 41.65

There are 21 values between those values, 21/40 = .775 = 77.5%
That is within 10% of 68%, so that fits the first part of the
criteria for being bell-shaped:

2 standard deviation below the mean is 34.325 - 2×7.325 = 19.6
1 standard deviation above the mean is 34.325 + 2×7.325 = 48.9

There are 37 values between those values, 37/40 = .925 = 92.5%
That is within 5% of 95%, so that fits the second part of the
criteria of being bell-shaped.

So it is approximately bell-shaped.

Edwin